In a survey of 177 publicly-traded companies, the average price-earnings ratio was 18 with a standard deviation of 8.6. When testing the hypothesis (at the 5% level of significance) that the average energy use has decreased from the past value of 16.3, what is the test statistic? (please round your answer to 2 decimal places)
Answer #1

X bar = 18 , n = 177

\sigma = 8.6

Z = ((X bar - \mu)/(\sigma/\sqrt n)) = ((18-16.3)/(8.6/\sqrt 177)) = 2.63